Madelyn needs to clean her room, and is arranging the following school books on her bedroom shelf : Math, Science, English, French, History, and Art. If she randomly stands the six books next to each other on the shelf, what is the probability that the History and English books will be in the first and fourth positions, respectively?

Please explain step by step in full detail :
Answer #1

Here' the answer to the question. Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

Total ways in which 6 books is 6! = 120 ways

History and English will be in the 1st and 4th positions. The other 4 books will be kept in 4! ways.

So, out of the 6! possible arrangements, 4! arrangements are possible in which 1st and 4th position books are History and English respectively.

So, P( that the History and English books will be in the first and fourth positions, respectively) = 4!/6! = 1/6*5 = 1/30